Additive Schwarz Algorithms for the h- and p-Versions of the Boundary Element Method

E.P.Stephan IfAM, Universitaet Hannover, Germany

Abstract

We study additive Schwarz algorithms for the h- and p- version boundary element method. Both weakly singular and hypersingular integral equations of the first kind are considered. We show that the condition number of the additive Schwarz operator is bounded for the h-version and grows at most as (log p)^{3} for the p-version. Numerical experiments are also given.